Your beginning salary is$70,000.Youdeposit12% at the end of each year in a savings account that earns 3% interest. Your salary increases by 2% per year. What value does your savings book show after 40 years?

please solve this question  on Excel sheet and showing the formula
Answer #1
Year Salary Deposit amount
1 70000 8400
2 71400.00 8568.00
3 72828.00 8739.36
4 74284.56 8914.15
5 75770.25 9092.43
6 77285.66 9274.28
7 78831.37 9459.76
8 80408.00 9648.96
9 82016.16 9841.94
10 83656.48 10038.78
11 85329.61 10239.55
12 87036.20 10444.34
13 88776.93 10653.23
14 90552.46 10866.30
15 92363.51 11083.62
16 94210.78 11305.29
17 96095.00 11531.40
18 98016.90 11762.03
19 99977.24 11997.27
20 101976.78 12237.21
21 104016.32 12481.96
22 106096.64 12731.60
23 108218.58 12986.23
24 110382.95 13245.95
25 112590.61 13510.87
26 114842.42 13781.09
27 117139.27 14056.71
28 119482.05 14337.85
29 121871.69 14624.60
30 124309.13 14917.10
31 126795.31 15215.44
32 129331.22 15519.75
33 131917.84 15830.14
34 134556.20 16146.74
35 137247.32 16469.68
36 139992.27 16799.07
37 142792.11 17135.05
38 145647.96 17477.75
39 148560.92 17827.31
40 151532.13 18183.86
NPV 2,71,412.68
Future worth 8,85,358.43

Showing formula in excel

Year Salary Deposit amount
1 70000 =B2*0.12
2 =B2*1.02 =B3*0.12
3 =B3*1.02 =B4*0.12
4 =B4*1.02 =B5*0.12
5 =B5*1.02 =B6*0.12
6 =B6*1.02 =B7*0.12
7 =B7*1.02 =B8*0.12
8 =B8*1.02 =B9*0.12
9 =B9*1.02 =B10*0.12
10 =B10*1.02 =B11*0.12
11 =B11*1.02 =B12*0.12
12 =B12*1.02 =B13*0.12
13 =B13*1.02 =B14*0.12
14 =B14*1.02 =B15*0.12
15 =B15*1.02 =B16*0.12
16 =B16*1.02 =B17*0.12
17 =B17*1.02 =B18*0.12
18 =B18*1.02 =B19*0.12
19 =B19*1.02 =B20*0.12
20 =B20*1.02 =B21*0.12
21 =B21*1.02 =B22*0.12
22 =B22*1.02 =B23*0.12
23 =B23*1.02 =B24*0.12
24 =B24*1.02 =B25*0.12
25 =B25*1.02 =B26*0.12
26 =B26*1.02 =B27*0.12
27 =B27*1.02 =B28*0.12
28 =B28*1.02 =B29*0.12
29 =B29*1.02 =B30*0.12
30 =B30*1.02 =B31*0.12
31 =B31*1.02 =B32*0.12
32 =B32*1.02 =B33*0.12
33 =B33*1.02 =B34*0.12
34 =B34*1.02 =B35*0.12
35 =B35*1.02 =B36*0.12
36 =B36*1.02 =B37*0.12
37 =B37*1.02 =B38*0.12
38 =B38*1.02 =B39*0.12
39 =B39*1.02 =B40*0.12
40 =B40*1.02 =B41*0.12
NPV =NPV(3%,C2:C41)
Future worth =FV(3%,40,,-C42)